Suppose you need to have $52,938.00 in an account 18.00 years from today and that the account pays 15.00%. How much do you have to deposit into the account 8.00 years from today?

Suppose you deposit $1,773.00 into and account 4.00 years from today into an account that earns 14.00%. How much will the account be worth 18.00 years from today?

Assume the nominal rate of return is 9.93% and the real rate is 2.96%. Find the inflation rate of return using the exact formula
Answer #1

The solution to the given question is provided below for your reference

1.
=52938/1.15^(18-8)=13085.46397

2.
=1773*1.14^(18-4)=11101.37196

3.
=1.0993/1.0296-1=6.7696193%